### v1

# Phase space maximal entropy random walk: Langevin-like ensemble of physical trajectories

Abstract
As written by statistician George Box ”All models are wrong, but some are useful”, standard diffusion derivation or Feynman path ensembles use nonphysical nowhere differentiable trajectories of infinite kinetic energy - what seems wrong, bringing question of differences if doing it more right this article is focused on. To consider ensembles of more physical trajectories, we can work in $(x,v)$ phase space like in Langevin equation with velocity controlling spatial steps, here also controlled with spatial potential $V(x)$. There will be discussed and compared 4 approaches to predict stationary probability distributions: using Boltzmann ensemble of points in space (GRW - generic random walk) or in phase space (psGRW), and analogously Boltzmann ensemble of paths in space (MERW - maximal entropy random walk) and in phase space (psMERW). They have qualitatively different predictions, hopefully allowing to decide the most appropriate for various settings by distinguishing them experimentally. Path ensembles have much stronger Anderson-like localization exactly as in quantum ground state, proposed novel in phase space has additionally increased density of velocity as in potential gradient, decreased toward barrier. While MERW is thermodynamically in agreement with quantum mechanics, psMERW suggests slight corrections from consideration of more physical trajectories, which seem natural - reduced velocities toward close barriers, increased down potential gradients.
Keywords: diffusion, phase space, Langevin equation, Schrödinger equation, maximal entropy random walk
I Introduction
In standard derivation of diffusion equations, or in Feynman path integrals [1] we consider steps in space, for continuous limit using $\epsilon\propto\delta^{2}\to 0$ for $\epsilon$ being time step and $\delta$ being spatial step, what means velocity and $mv^{2}/2$ kinetic energy going to infinity ($\delta/\epsilon\to\infty$), nowhere differentiable trajectories.
While such trajectories are clearly nonphysical, a basic question is if it is in an issue? It is mathematically more difficult, but doable to use ensembles of more physical trajectories - the question is if it could lead to a better agreement with experiment? The main purpose of this article is to start asking these basic but very difficult questions, by comparing 4 different approaches for a given spatial potential $V(x)$, with example predicted stationary probability distributions summarized in Fig. 1.
Figure 1: Examples of 4 types of predictions for 3 basic scenarios: 1D harmonic potential, 1D infinite well, and radial 3D $V\propto-1/r$ point electric or gravitational potential. Top: predicted $\rho(x)$ spatial stationary probability densities - for GRW and psGRW it is Boltzmann distribution in space. For MERW and QM it is Boltzmann path ensemble in space - $\rho\propto\psi^{2}$ for satisfying (Schrödinger) eigenequation $E\psi(x)=(-\Delta_{x}+V(x))$ for minimal energy $E$. For proposed psMERW it is Boltzmann path ensemble in phase space - as for MERW but for $E\psi(x,v)=(-\Delta_{v}-2v\cdot\nabla_{x}+V(x)+\|v\|^{2})\,\psi(x,v)$ eigenequation (found with Wolfram Mathematica ”NDEigensystem”). Center: stationary probability density in phase space for psGRW (upper) and psMERW (lower). Below are its chosen cross-sections for positions shown with colors. For infinite potential well, near barriers only opposite velocities should be allowed - what is seen here only for psMERW. For harmonic potential we see increased density of velocity down the potential gradient.
To consider ensembles of more physical trajectories, as in Langevin equation [2] we can go to $(x,v)$ phase space - include finite velocity $v$ in evolved state, which evolves randomly and controls evolution in space $(x_{t+1}-x_{t}\propto v_{t})$. To get various behaviors which could be compared experimentally, there is included general spatial potential $V(x)$.
Another basic crucial question is if we should base on maximizing local entropy production GRW (generic random walk) with walker considering single steps, what leads to Boltzmann ensemble in space, or maybe better maximizing mean entropy production MERW (maximal entropy random walk) using Boltzmann ensemble of entire paths - like random walk along Ising sequence, or ”Wick-rotated” Feynman path ensembles, leading to the same stationary probability distribution as quantum ground state, with much stronger Anderson-like localization property [3].
This much stronger QM/MERW localization is crucial especially for electrons in semiconductor - standard diffusion would predict nearly uniform stationary electron probability distribution for such lattice of usually two types of atoms, which should flow if attaching external potential. In contrast, in QM/MERW and experiments these electrons are strongly localized, visualized e.g. in [4], what prevents conductance in such semiconductor. MERW allows for working diffusions model of e.g. diode as semiconductor p-n junction [5].
Neutrons are another objects confirmed to have quantum ground state stationary probability distribution [6] as predicted by both MERW and QM. It brings a difficult general question which approach should be used in which case, e.g. in diffusion of molecules, solitons, or dust halos in astronomical settings.
Figure 2: Three basic philosophies to predict stationary probability distributions. Left: in stochastic picture we guess stochastic propagator, usually maximizing entropy separately for each step (GRW), what leads to Boltzmann ensemble of positions in space. Right: in ergodic picture we assume we fully control deterministic evolution. Center: in contrast, in thermodynamics, statistical physics we are not so certain we know such details. Instead, we assume the safest statistical model accordingly to Jaynes principle: maximizing entropy. For each point we consider Boltzmann ensemble of possible paths from it, their statics give MERW stochastic propagator. It is nonlocal (depends on the entire situation), but effective model - used only by us to predict the most likely distributions, not used by the object itself - governed by some complex unknown evolution. In this article we expand MERW philosophy to phase space, using ensemble of more physical trajectories.
While MERW might be appropriate for frequently interacting objects, e.g. for rarely interacting sparse dust halo it seems more connivent to consider MERW in phase space. Combining them was one of motivations for this article, also to compare predictions of all these models, which hopefully could be experimentally distinguished in some future.
Beside Langevin equation, related approaches are e.g. phase-space formulation of quantum mechanics [7], or Feynman path ensemble in phase space [8]. However, they don’t actually use velocities to choose spatial step, leading to different equations.
While for psGRW the stationary densities as just products of Boltzmann distributions for space and velocity, for psMERW they interact - leading to looking obvious physical properties as we can see in Fig. 1: in gradient of potential there is increased density of velocity down this gradient, near potential barriers there is decreased density of velocity toward it.
Unfortunately the formulas become more complicated, requiring to solve functional eigenequations, for which analytical solutions could be found rather only for very simple cases. Fortunately there are available tools to do it numerically, like the used NDEigensystem function of Wolfram Mathematica.
Section II introduces to GRW and MERW philosophies, Section III contains their proposed extension to phase space. This is initial version of article, hopefully stimulating further research in this direction, especially toward experimental distinguishing of the discussed 4 approaches.
Figure 3: Gathered GRW, MERW formulas for graph defined by matrix $M$.
II Generic and maximal entropy random walk
This section briefly introduces to GRW/MERW [^1] MERW introduction: https://community.wolfram.com/groups/-/m/t/2924355 random walks summarized in Fig. 2, 3, for details see e.g. [3, 9].
While further we will use lattice for continuous limit, let us start with general random walk on a graph given by adjacency matrix $M$: $M_{ij}=0$ if there is no edge between $i$ and $j$, otherwise it is 1 or generally has some weight, for Boltzmann ensemble useful to parameterize $M_{ij}=\exp(-\beta E_{ij})$ for $E_{ij}$ being step energy and thermodynamical $\beta=1/k_{B}T$.
For such given $M$ matrix we would like to find stochastic matrix of transition probabilities for Markov process:
$S_{ij}=\textrm{Pr}(X_{t}=j|X_{t-1}=i)\quad\textrm{satisfying}\quad\sum_{j}S_{ij}=1$
and using only allowed transitions: $M_{ij}=0\Rightarrow S_{ij}=0$, what is equivalent with assigning $E_{ij}=\infty$ energy to this transition.
A basic approach to choose $S$, referred as generic random walk (GRW), is assuming uniform/Boltzmann ensemble among single possible steps, this way maximizing local entropy for this step (minus mean energy for Boltzmann):
$S^{GRM}_{ij}=\frac{M_{ij}}{d_{i}}\quad\textrm{for}\quad d_{i}=\sum_{j}M_{ij}$ (1)
For symmetric $M=M^{T}$ it leads to stationary probability distribution $\rho S=\rho$ being just $\rho_{i}^{GRW}=d_{i}/\sum_{j}d_{j}$.
II-A Maximal entropy random walk (MERW)
In contrast, in MERW we maximize mean entropy production (minus mean energy for Boltzmann), what is equivalent to uniform/Boltzmann ensemble among infinite paths.
The basic tool to calculate its stationary distribution and propagator is noticing that $M^{l}$ matrix power for $M_{ij}=\exp(-\beta E_{ij})$, combinatorially contains sum of Boltzmann path ensemble:
$(M^{l})_{ij}=\sum_{\gamma:\gamma_{0}=i,\gamma_{l}=j}\exp(-\beta E(\gamma))\qquad\textrm{for}\ \ E(\gamma)=\sum_{i=1}^{l}E_{\gamma_{i-1}\gamma_{i}}$ (2)
We are interested in $l\to\infty$ limit, for which using natural assumption of connected and acyclic graph, the Frobenius-Perron theorem says there is a single dominant eigenvalue $\lambda$, allowing to use the dominant eigenvector $\psi$ in the limit, amplitude as result of propagator from past or future infinity:
$M^{l}\propto\lambda^{l}\psi\psi^{T}\equiv\lambda^{l}|\psi\rangle\langle\psi|\qquad\textrm{for}\quad\lambda\to\infty$ (3)
We can use both (2) and (3) to find propagator and stationary probability distribution (for any $j,k$) as in Fig. 4:
$S_{ij}^{MERW}=\frac{\textrm{Pr}(ij)}{\textrm{Pr}(i)}=\lim_{l\to\infty}\frac{M_{ij}(M^{l})_{jk}}{(M^{l+1})_{ik}}=\frac{M_{ij}}{\lambda}\frac{\psi_{j}}{\psi_{i}}$ (4)
$\rho^{MERW}_{i}\propto\lim_{\l\to\infty}(M^{l})_{ji}(M^{l})_{ik}\propto(\psi_{i})^{2}$ (5)
Figure 4: Diagrams for derivation of MERW formulas from path ensembles, exploiting that $(M^{l})_{ij}$ combinatorially contains summation over uniform/Boltzmann ensemble of length $l$ paths from $i$ to $j$. To get stochastic propagator $S^{MERW}_{ij}$, we can consider length $l$ paths from vertex $i$, and look at distribution of their first step for $l\to\infty$ limit. To get stationary probability distribution we can multiply such two propagators: $\rho^{MERW}_{i}\propto\lim_{l\to\infty}(M^{l})_{ji}(M^{l})_{ik}$.
II-B Lattice and continuous limit to Schrödinger equiation
Let us discretize space $\mathbb{R}^{D}\ni x=\delta\tilde{x}$ for $\tilde{x}\in\mathbb{Z}^{D}$ and time $\mathbb{R}\ni t=\epsilon\tilde{t}$ for $\tilde{t}\in\mathbb{Z}$. For continuous limit $\epsilon,\delta\to 0$. Also disretize spatial potential: $V_{\tilde{x}}\equiv V(\delta\tilde{x})$. For simplicity let us work in $D=1$, with mentioned generalization for $D>1$.
II-B1 GRW leading to Boltzmann space ensemble
Allowing only jumps up to the nearest neighbors: step $s\in\{-1,0,1\}$, we can choose $M$ matrix as symmetric and using potential:
$M_{\tilde{x},\tilde{x}+s}=\exp(-\beta((V_{\tilde{x}}+V_{\tilde{x}+s})/2)\quad\textrm{for }s\in\{-1,0,1\}$ (6)
as it symmetric, GRW stationary probability distribution is $\propto M_{\tilde{x},\tilde{x}-1}+M_{\tilde{x},\tilde{x}}+M_{\tilde{x},\tilde{x}+1}$. Assuming continuous potential $V(x)$, in continuous limit we get just Boltzmann ensemble in space:
$\rho^{GRW}(x)\propto\exp(-\beta V(x))$ (7)
II-B2 MERW leading to Schrödinger equiation
[…]
$M_{\tilde{x},\tilde{x}+s}=\exp(-\beta\epsilon((V_{\tilde{x}}+V_{\tilde{x}+s})/2)$
For MERW we first need to find the dominant eigenvector of adjacency matrix $\lambda\psi=M\psi$ for $M$ given by (6):
$\lambda\psi_{\tilde{x}}=(M\psi)_{\tilde{x}}=\sum_{s\in\{-1,0,1\}}e^{-\beta\epsilon(V_{\tilde{x}}+V_{\tilde{x}+s})/2}\ \psi_{\tilde{x}+s}$ (8)
Taking $\exp(-\epsilon)\approx 1-\epsilon$ first order expansion and assuming continuous potential $V(x)$, for continuous limit we can approximate the above with:
$\lambda\psi_{\tilde{x}}\approx\psi_{\tilde{x}-1}+\psi_{\tilde{x}}+\psi_{\tilde{x}+1}-3\beta\epsilon V_{\tilde{x}}$
Subtracting $3\psi_{\tilde{x}}$ from both sides, and multiplying by $-1/3\beta\epsilon$:
$\frac{3-\lambda}{3\beta\epsilon}\psi_{\tilde{x}}\approx-\frac{1}{3\beta}\frac{\psi_{\tilde{x}-1}-2\psi_{\tilde{x}}+\psi_{\tilde{x}+1}}{\epsilon}+V_{\tilde{x}}\psi_{\tilde{x}}$
Due to change of sign, eigenvalue maximization becomes minimization of $(3-\lambda)/(3\beta\epsilon)$, which in continuous limit should converge to some energy $E$. Assuming $\epsilon=\delta^{2}$ the difference term will lead to laplacian, hence in continuous limit we can write the eigenequation as stationary Schrödinger equitation searching for the lowest energy ground state:
$E\psi=-C\Delta\psi+V\psi\qquad\rho^{MERW}(x)\propto(\psi(x))^{2}$ (9)
for $C=1/3\beta$ in 1D or generally $1/(2D+1)$ in dimension $D$. We can choose parameters to make it $C=\hbar^{2}/2m$ as in standard Schrödinger equitation, what is explored e.g. in  [9] or in [10] finding also some agreement toward Dirac equation.
III Random walks in phase space
While standard random walks or Feynman path ensembles use nowhere differentiable paths of infinite kinetic energy, here we would like use more physical paths by going to phase space $(x,v)$: with random change of (finite) velocity $v$, which defines deterministic change of position $x$.
Like in popular Langevin approach [2]:
$m\frac{dx}{dt}=v\qquad m\frac{dv}{dt}=-\Gamma v+\eta$
for $\Gamma$ describing damping (neglected in current version). It has no potential, hence independent finite variance infinitesimal steps would lead to Gaussian distribution for velocities.
Here we would like to include energy - both kinetic $mv^{2}/2$ and some general position dependant $V(x)$:
$\bar{V}(x,v)=V(x)+m\|v\|^{2}/2$ (10)
Like discussed previously, performing continuous limit of $M_{(\tilde{x},\tilde{v}),(\tilde{x}+\tilde{v},\tilde{v}\pm 1)}\sim\exp(-\bar{V}(x,v))$ for phase space GRW (psGRW), the stationary probability distribution would be just:
$\rho^{psGRW}(x,v)\propto\exp(-\beta(V(x)+mv^{2}/2))$ (11)
III-A MERW continuous limit in phase space
Let us now discretize space $\mathbb{R}^{D}\ni x=\delta\tilde{x}$ for $\tilde{x}\in\mathbb{Z}^{D}$ and time $\mathbb{R}\ni t=\epsilon\tilde{t}$ for $\tilde{t}\in\mathbb{Z}$. For continuous limit we will take $\epsilon,\delta\to 0$. Discretized $\tilde{v}\in\mathbb{Z}^{D}$ velocity for $\tilde{x}\to\tilde{x}+\tilde{v}$ transition in single step corresponds to $v=\frac{\delta}{\epsilon}\tilde{v}\in\mathbb{R}^{D}$ real velocity. As previously we calculate for $D=1$, then mention for general $D$.
Denote $\psi_{\tilde{x},\tilde{v}}=\psi(\delta\tilde{x},\frac{\delta}{\epsilon}\tilde{v})=\psi(x,v)$ as discretization of continuous eigenfunction in phase space. Analogously for potential: $\bar{V}_{\tilde{x},\tilde{v}}=\bar{V}(\delta\tilde{x},\frac{\delta}{\epsilon}\tilde{v})=\bar{V}(x,v)$.
In one step $\tilde{x}\to\tilde{x}+\tilde{v}$, and there is a random transition of $\tilde{v}$, we can assume at most ±$\pm 1$. To include damping in future, we could add velocity reduction in below step. In 1D the MERW eigenequation after neglecting higher order terms becomes:
$\lambda\psi_{\tilde{x}-\tilde{v},\tilde{v}}=(M\psi)_{\tilde{x}-\tilde{v},\tilde{v}}=\sum_{s\in\{-1,0,1\}}e^{-\epsilon\beta V_{\tilde{x},\tilde{v}+s}}\psi_{\tilde{x},\tilde{v}+s}\approx$ (12)
$\approx\psi_{\tilde{x},\tilde{v}-1}+\psi_{\tilde{x},\tilde{v}}+\psi_{\tilde{x},\tilde{v}+1}-3\beta\epsilon V_{\tilde{x},\tilde{v}}\psi_{\tilde{x},\tilde{v}}$
Approximating using derivative $\nabla_{x}\equiv\partial/\partial x$:
$\psi_{\tilde{x}-\tilde{v},\tilde{v}}=\psi(\delta(\tilde{x}-\tilde{v}),v)=\psi(x-\epsilon v,v)\approx(\psi-\epsilon v\nabla_{x}\psi)_{\tilde{x},\tilde{v}}$
and substituting it to eigenequation (12) we get:
$\lambda(\psi-\epsilon v\nabla_{x}\psi)_{\tilde{x},\tilde{v}}\approx\psi_{\tilde{x},\tilde{v}-1}+\psi_{\tilde{x},\tilde{v}}+\psi_{\tilde{x},\tilde{v}+1}-3\beta\epsilon V_{\tilde{x},\tilde{v}}\psi_{\tilde{x},\tilde{v}}$
Now subtract $3\psi_{\tilde{x},\tilde{v}}$ from both sides and multiply by $-1/(3\beta(\delta/\epsilon)^{2})$, getting:
$\frac{3-\lambda}{3\beta(\delta/\epsilon)^{2}}\psi_{\tilde{x},\tilde{v}}+v\frac{\lambda}{3\beta}\frac{\epsilon}{(\delta/\epsilon)^{2}}(\nabla_{x}\psi)_{\tilde{x},\tilde{v}}\approx$
$\approx-\frac{1}{3\beta}\frac{\psi_{\tilde{x},\tilde{v}-1}-2\psi_{\tilde{x},\tilde{v}}+\psi_{\tilde{x},\tilde{v}+1}}{(\delta/\epsilon)^{2}}+\frac{\epsilon}{(\delta/\epsilon)^{2}}V_{\tilde{x},\tilde{v}}\psi_{\tilde{x},\tilde{v}}$
Assuming $\delta/\epsilon\to 0$, the discrete laplacian tends to continuous $\Delta_{v}$ for velocity. As previously we would like to interpret the first coefficient as energy: $\frac{3-\lambda}{3\beta(\delta/\epsilon)^{2}}\to E$.
For $\nabla_{x}$ gradient term we can choose $(\lambda\to 3)$:
$\frac{\lambda}{3\beta}\frac{\epsilon}{(\delta/\epsilon)^{2}}\to\xi\qquad\textrm{for example}\quad\epsilon=\sqrt[3]{\beta\xi\,\delta^{2}}$ (13)
or use a sequence $\xi\to 0$ for $\xi=0$ in continuous case.
Finally in the continuous limit we get stationary phase space Schrödinger equation for $\psi\equiv\psi(x,v)$:
$E\psi=-C\Delta_{v}\psi-\xi v\cdot\nabla_{x}\psi+(V+m\|v\|^{2}/2)\psi$ (14)
where $C=\frac{1}{3\beta}$, or generally $C=\frac{1}{(2D+1)\beta}$ in dimension $D$, is some constant which can be directly chosen, e.g. dependent on $\hbar$ or temperature. Analogously for $\xi$, down to $\xi\to 0$ limit making position and velocity independent.
To find the stationary probability distribution $\rho(x,v)$, we need to solve eigenequation (14) minimizing energy $E$ (higher energy contributions have exponential decay), then $\rho(x,v)\propto\psi(x,v)^{2}$.
While for QM/MERW there are well known analytical solutions of such stationary Schrödinger equation e.g. for harmonic potential, infinite well, point $-1/r$ potential, their phase space version are more complicated and Wolfram Mathematica DEigensystem was not able to find analytical solutions, hence for their example solution in Fig. 1 there was used numerical NDEigensystem, also required for more complex settings.
Function NDEigensystem uses finite element method, and already had various issues in the discussed basic cases - resolved e.g. by manually specifying Dirchlet conditions for infinite well, or for point potential: solving for $\psi(x,v)\sim\exp(-v^{2}-x)\phi(x,v)$ type substitution explicitly assuming asymptotic behavior. Alternative approach is approximation by solving in a finite orthogonal basis of functions (removing higher modes), especially Gaussian distribution times Hermite polynomial for velocities (ground state already gets such higher modes also for velocity). Numerical methods for psMERW will require further work, especially to include interaction between multiple particles.
IV Conclusions and further work
While popularly considered path ensembles usually use infinite kinetic energy paths, it is mathematically more difficult but doable to use ensembles of more physical paths instead by going to phase space. This article discusses some basics with included spatial potential to bring attention, hopefully leading to some attempts to experimentally determine the most appropriate ones for various physical settings.
This is initial article opening various directions for further work, e.g.:
  • Search for possibilities of experimental determination which of the discussed approaches is the most appropriate, especially using stationary probability distributions of positions, maybe also velocities e.g. from redshifts - in various settings from microscopic e.g. molecules, solitons to astronomical e.g. dust halos.
  • Find parameters, understand their dependence, universality.
  • Mathematical improvements, especially of numerical approaches, also search for analytical formulas for basic cases, maybe include damping like in Langevin equation, larger velocity changes through collisions, include relativistic corrections, etc.
  • While for simplicity we have discussed only stationary situation, it seems worth to also consider dynamics - both straightforward evolution in fixed potential like for electron conductance [5], but also much more difficult case of varying potential - with basics discussed in [9].
  • While the discussed analysis was for a single walker, in practice we usually have multiple interacting - what should be finally included. One way is through mean-field treatment assuming the remaining have the same distribution, used e.g. for MERW conductance model [5] to include potential contributions from the remaining electrons. For psMERW it could be useful e.g. for astronomical dust halos to include their gravitational self-interaction.
  • While we have focused on continuous limit of infinitesimal lattice constants, finite lattice can be considered e.g. for conductance models imagining electrons jumping between atoms in a lattice. Also, for MERW Darwin term was recently derived [10] as correction from use of finite lattice. Especially the latter suggests to also consider psMERW in some finite lattice, and closely look at corrections it brings.